C12H22O11 + H2O = C6H12O6 + C6H12O6

How many mol of glucose can be produced from 98 mol of sucrose?
1
Expert's answer
2020-11-20T08:08:48-0500

C12H22O11 + H2O = C6H12O6 + C6H12O6

"\frac{98}{1}=\frac{x}{1}"

x=98 mol